OpenStudy (anonymous):

Suppose in a proof of a summation formula 1+5+25+...+5^k -1=1/4 (5^k -1) by mathematical induction, you show the formula valid for L=1 and assume that it is valid for L=k. What is the next equation in the induction step of this proof? Please help! Will give best answer!

OpenStudy (anonymous):

Is it a written response or multiple choice?

OpenStudy (anonymous):

BTW i like ur profile pic ;)

OpenStudy (anonymous):

multiple choice. im writing them out now

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Where are the answers? :?

OpenStudy (anonymous):

a 1+5+25+...+5^k-1 +5^k+1-1 = 1/4(5^k -1) +1/4(5^k+1 -1) b 1+5+25+...+5^k +5^k+1 = 1/4(5^k -1)+5^k+1-1

OpenStudy (anonymous):

are there more?

OpenStudy (anonymous):

yes they just take forever to write so hold on

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

c 1+5+25+...+5^k-1 +5^k+1-1 =1/4 (5^k -1) + 5^k+1-1 d 1+5+25+...+5^k-1 +5^k+1-1 = 1/4 (5^k+1 -1) + 5^k+1-1

OpenStudy (anonymous):

ok the anser is c. i think im not so sure

OpenStudy (anonymous):

plz become my fan ok

OpenStudy (anonymous):

why do you think that?

OpenStudy (anonymous):

well because i just think so...

OpenStudy (anonymous):

im not going to become a fan, if you don't actually help me